OpenStudy (anonymous):

Evaluate: anti-derivative x dx/sqrt(x^2+1)

OpenStudy (anonymous):

I get u= x^2+1 du = 1 dx

OpenStudy (anonymous):

put x^2 +1 =t diff w.r.t. x , we get 2x dx= dt

OpenStudy (anonymous):

i.e. x dx= dt/2

OpenStudy (anonymous):

u could even substitute sqrt(x^2+1)=t than can work too

OpenStudy (amistre64):

tan(t) = x ; dt sec^2 = dx; t = tan^-1(x)..lets start :) x dx tan(t) sec^2(t) dt -------- --> --------------- (x^2 +1) tan^2 +1 <-- this is eqaul to sec^2

OpenStudy (amistre64):

[S] tan(t) dt

OpenStudy (anonymous):

(x^2 + 1)^1/2 + c

OpenStudy (amistre64):

yeah, it helps if you read the sqrt part ;).....

OpenStudy (amistre64):

...if you ever need to do the one I was working on....lol

OpenStudy (amistre64):

when I re did it with the approriate sqrt.... I got\[\int\limits_{}\frac{\tan(t)\sec^2(t)}{\sec(t)}dt \rightarrow \int\limits_{} \sec(t)\tan(t)dt \rightarrow \sec(t)\]

OpenStudy (amistre64):

\[\sec(t) = \sqrt{x^2+1}\]

OpenStudy (amistre64):

+C lol

OpenStudy (anonymous):

I started a new problem

OpenStudy (amistre64):

yay!! .... I was just trying to wrap this one up in me brain :)